How much heat is transferred in the process of compleay melting a 1.60kg block of ice starting at -15 degrees celcious
1
Expert's answer
2019-09-30T04:53:17-0400

The total amount of heat is the sum of heat necessary to heat the block of ice to melting point ("0[^\circ {\rm{C]}}") and heat necessary to melt this block of ice


"q = {q_{{\rm{heating}}}} + {q_{{\rm{melting}}}} = cm\Delta T + \lambda m"

where "c \approx 2100[\frac{{\rm{J}}}{{{\rm{kg}} \cdot {\rm{^\circ C}}}}]" is the specific heat capacity of ice (we neglecting its temperature dependence) and "\lambda \approx 330 \cdot {10^3}[\frac{{\rm{J}}}{{{\rm{kg}}}}]" is the heat of fusion of ice. Now let's perform calculations


"q = 2100[\frac{{\rm{J}}}{{{\rm{kg}} \cdot {\rm{^\circ C}}}}] \cdot 1.6[{\rm{kg}}] \cdot 15[^\circ {\rm{C]}} + 330 \cdot {10^3}[\frac{{\rm{J}}}{{{\rm{kg}}}}] \cdot 1.6[{\rm{kg}}] = 578400[{\rm{J}}]"

This is the answer.